OpenStudy (brucebaner):

Which is an equation in point-slope form for the given point and slope? Point:(8,-4) ; Slope: -5/6 A y+4=-5/6(x+8) B y-4=-5/6(x+8) C y+4=-5/6(x-8) D y-4=5/8(x-8)

OpenStudy (aripotta):

so point-slope form is y - y1 = m(x - x1), where m is the slope. just plug in the info into the formula to get your answer

OpenStudy (brucebaner):

okay so its B

OpenStudy (aripotta):

y + 4 = -5/6(x -8)

OpenStudy (aripotta):

so the answer is c

OpenStudy (brucebaner):

thx so much

OpenStudy (aripotta):

the point is (8,-4) and when you subtract a negative number, it becomes addition

OpenStudy (aripotta):

so instead of y - (-4), it becomes y + 4

OpenStudy (aripotta):

make sense?